Use production function: Q = K ^0.42 L^ 0.5 with Wage = 10, Rent = 6, output Price = $22/unit. To the nearest 0.01, what is the optimal quantity of Labor?
Answer #1

SOLUTION Whe а ее W 10 : 4 22 Th o the Thus heseme pme labos O L& k 2 a L Ak ad dib ad This conditiem atates-at he ma matuthe The cedition Thars,the ㅠ 22( K042 10.5) A TI 22(0 42)05,6 0-42 0.5 22 0.5) k 0420.5 ㄥ lo 20 (2)一0-5% 0.5 924 K 6 -0.5% 0.5 9-24 0 5% 1-24 し0.5 2-2 Fronea (2) 0,42 0.5し0.5 4- lo n (3) and (t) 0-42 l o 24 11x9.24 - lo X b Ix9.2 0-1な loX6 lox 6 = 26,960-42 0-5 0-42 L | 26.96 lo L = 19.26 Thug the-spurn a 1- award labor